OpenStudy (anonymous):

The length of a rectangle is 2m more than the width. The area of the rectangle is 20 m^2. Find the dimensions of the rectangle, to the nearest tenth of a metre.

OpenStudy (anonymous):

The answer is 3.6 m by 5.6m... how?

OpenStudy (anonymous):

We know that for a rectangle, the area is length x width A = l*w We know that the length is 2 meters more than the width So the the length minus 2 in the width => l - 2 = w l = w + 2 Substitutin in the area formula: A = (w+2)(w) 20 = w(w+2) Solve for w using the quadratic equation, then using the w you can find l

OpenStudy (anonymous):

solve for w using the quadratic equation? so it's going ot be w^2 + 2w - 20 = 0

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ohhhh,

OpenStudy (anonymous):

Get it now? :)

OpenStudy (anonymous):

yup! Oh, but what's a product of -20 and a sum -20? o.o

OpenStudy (anonymous):

Product of -20, sum of 2... Do you know how to use the quadratic equation?

OpenStudy (anonymous):

sort of.. isn't it like factoring?

OpenStudy (anonymous):

Its like factoring, but when like integer numbers dont work to solve the equation http://www.oncalc.com/wp-content/uploads/2011/02/quadratic-equation-calculator.gif

OpenStudy (anonymous):

You see that equation before?

OpenStudy (anonymous):

ohhhh the quadratic formula?

OpenStudy (anonymous):

Yeah, sorry didnt know what you call it

OpenStudy (anonymous):

it's okay. But what about it? That's all I use after I've figured it's w^2 + 2w - 20 =0?

OpenStudy (anonymous):

Yup, plug it in to the quadratic formula

OpenStudy (anonymous):

ohh, thank you!

OpenStudy (anonymous):

Okay, well I got -2 +/- 2sqr.rt:21/2 ...

OpenStudy (anonymous):

punch that into a calculator and what do you get?

OpenStudy (anonymous):

so I don't have to do 2(-2 +/- sqr.rt:21/2?

OpenStudy (anonymous):

well what number do you get from that?

OpenStudy (anonymous):

Like evaluate the expression

OpenStudy (anonymous):

not the answer.

OpenStudy (anonymous):

(-2 + sqrt ( 4 - 4(1)(-20) )/ 2

OpenStudy (anonymous):

I got 3.58

OpenStudy (anonymous):

if you do (-2 - sqrt ( 4 - 4(1)(-20) )/ 2 , you get a negative number, and can omit it because there is no such thing as negaive lengths